OpenStudy (anonymous):

2(2v+3)+8v-6(v-1)+3v solve the equation

OpenStudy (anonymous):

Its not an equation, its an algebraic expression.

OpenStudy (anonymous):

What would you do ?

OpenStudy (anonymous):

I would: take (2v+3)=5vx2=10v=2v-6v=4v+3v this is what I got

OpenStudy (anonymous):

You can't sum up two things that doens't have the same expression. so 2v + 3= 2v +3 ≠ 5v

OpenStudy (anonymous):

2(2v+3)+8v-6(v-1)+3v = 4v + 6 + 8v - 6(v-1) +3v ... continue...

OpenStudy (anonymous):

it's Never too late.

OpenStudy (anonymous):

2(2v+3)+8v-6(v-1)+3v = 9v+12 Set v to any number, say, 13 and evaluate both sides of the above equation. 129 = 129

OpenStudy (anonymous):

2(2v+3)+8v-6(v-1)+3v = 4v + 6 + 8v - 6v + 6 + 3v = 9v + 12. By bad.